ok. i need help finding th surface area of a cylindrical can 15 cm high and with a diameter of 8 cm
Answer by checkley77(12844)

THE TOP & BOTTOM AREA=2*PIR^2=2*3.14*4^2=100.48 CM^2
THE SIDE AREA=2PIR*15=2*3.14*4*15=376.8 CM^2
100.48+376.8=477.28 CM^2 IS THE TOTAL SURFACE AREA.